# Universal Frequency Correlations and Recurrence Statistics of Complex Impedance Matrices

### v1

@@ 1 Introduction
Random Matrix Theory (RMT) based models have emerged as especially successful at predicting certain universal properties of wave chaotic systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. One such model, the Random Coupling Model (RCM), was developed to incorporate nonuniversal physical features that practical systems invariably have, such as coupling details of ports or antennas [20, 21] and short orbits (also known as direct processes) [22, 23, 24], with an RMT description of the wave propagation within the enclosure. The RCM models the system including its ports in terms of an impedance or admittance matrix that linearly relates the voltages and currents appearing at the system’s ports. From these matrices, combined with knowledge of the external transmission systems feeding the ports, it is possible to evaluate the system scattering matrix, which is often of interest.
The RCM has been applied to understanding the statistical electromagnetic properties of complex enclosed systems for decades [25, 26, 27], and has also been empirically verified to describe acoustic scattering [28]. Predictions of the RCM have been compared to experimental scattering and impedance matrices [25, 29, 23, 30, 31, 32], admittance matrices [33], conductance [34], fading statistics [35], complex time delay statistics [36, 37], and scattering singularity statistics [38].
One of the applications of the RCM is identifying the underlying universal fluctuations of impedance from experimental ensemble data, which is done by normalizing the measured cavity impedance matrix by the radiation impedance specific to the ports or antennas used in the measurement. The resulting “normalized impedance” matrix $Z$ is then dimensionless and independent of coupling details. In this paper, we examine the RCM normalized impedance matrix obtained from experimental data, and identify characteristics of the two point frequency correlation functions of the real (resistance $R$) and imaginary (reactance $X$) components of the impedance matrix elements.
@@
We then apply the knowledge of the correlation functions to address the question: If a port of the system presents a particular value of the complex impedance $Z^{\text{t}}$ at a frequency $f$, what is the average frequency interval $\Lambda$ before the same value of impedance $Z^{\text{t}}$, within some specified tolerance, is seen again? An example is shown in Fig. 1 where we show the path in the complex plane traced by an off-diagonal element of the RCM normalized impedance matrix, $Z_{12}$, over a short frequency range. The arbitrarily chosen target impedance value is $Z_{12}^{\text{t}}=0.082+0.114i$ marked by the black cross, and the small gray circle around it represents a tolerance threshold of $\epsilon=0.015$. The color bar represents the frequency difference $\delta=f_{a}-f_{b}$ (where $f_{b}$ is the first frequency at which $Z_{12}$ is within the gray circle) normalized by a characteristic scale corresponding to the typical $Q$-widths of the modes $\gamma=2\alpha\Delta=0.66~\text{MHz}$, where $\alpha$ is the dimensionless RCM loss parameter and $\Delta$ is the mean frequency spacing of the modes.
@@
Empirically we find that the expected frequency interval $\Lambda$ between repetitions of a target impedance depends on the mean mode spacing $\Delta$ of the system, the degree of absorption $\alpha$ of the system, the value of the desired target impedance $Z^{\text{t}}$, and the tolerance $\epsilon$ on the target value. We claim the scaling of $\Lambda$ to be interesting for several reasons. $\Lambda$ is a new statistical observable for chaotic scattering directly analogous to a recurrence time in dynamical chaotic systems, and has practical relevance for system engineering. Our scaling law provides a predictive tool for how often a desired impedance condition occurs naturally in frequency for a given cavity or resonator. This is also the first step towards a more difficult question, which would be looking for repetitions not of individual impedance elements, but of special conditions on the entire $Z$ matrix, such as impedance matrix Exceptional Points. These have previously been shown to be identical to scattering matrix Exceptional Points [39], so the impedance matrix may serve as a useful platform to study these and other wave scattering singularities.
@@
In this manuscript, we first present the theoretical predictions for the correlation functions and repetition interval $\Lambda$ in Sec. 2. The experimental systems used are introduced in Sec. 3, alongside an explanation of how we apply the Random Coupling Model to the raw data and the statistical methods used in the analysis. The results are shown and interpreted in Sec. 4, and we conclude our work and discuss future research in Sec. 5.
@@ 2 Theory
General forms of two point frequency correlation functions for the fluctuations of the impedance and scattering matrices were first proposed by Fyodorov, Savin, and Sommers [40, 41]. Various scattering matrix correlation functions have been studied since, often as a way to probe what is known as the elastic enhancement factor [33, 42, 43, 31, 14, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. In this paper, we consider the impedance two point autocorrelation functions, which have not been experimentally investigated before, to our knowledge.
Under the limiting assumptions of ideal coupling (which we achieve through the RCM normalization process), reciprocity (Dyson symmetry class $\beta=1$), and large loss ($\alpha\gg 1$), we arrive at the reduced expressions for the impedance two point frequency correlation functions for the elements of the RCM normalized impedance matrix with condensed notation:
@@
In these equations, $\delta=|f_{a}-f_{b}|$ is the difference in frequency between the two points, $\gamma=2\alpha\Delta$ is the characteristic frequency interval specific to the system, and $R^{\text{RMS}}=\sqrt{\langle(R_{pp^{\prime}})^{2}\rangle}$ is defined as the standard deviation of the fluctuation of the real component of an off-diagonal element of the RCM normalized impedance matrix. Note that in the case of large $\alpha$, $R^{\text{RMS}}$ is known to depend only on $\alpha$ by the formula $R^{\text{RMS}}=(\pi\alpha)^{-1/2}$ [25, 20].
In Fig. 2 we compare Eq. 1 to the real part of the impedance correlation function given in Eqs. [10a-b] of Ref. [40] with $\beta=1$ (reciprocity), all $\kappa=1$, (ideal coupling), and three different $\alpha$ values. We see that for large $\alpha$, our equation and the expression derived by Fyodorov, Savin, and Sommers for the two point frequency correlation function of the real part of the diagonal impedance elements are equivalent.
@@
Figure 2: Two point frequency correlation function of the real part of a diagonal impedance element $\langle\tilde{R}(f_{a})\tilde{R}(f_{b})\rangle_{C}$ normalized by the zero separation correlation $\langle(\tilde{R}_{pp})^{2}\rangle$. Comparing prediction of Eq. 1 illustrated by the dashed black curve, and expression derived by Fyodorov, Savin, and Sommers in Ref. [40], represented by the colored curves corresponding to different choices of loss parameter $\alpha$. The green curve which has $\alpha=10$ is indistinguishable from the dashed black curve.
@@
The correlation functions for the diagonal elements come out to be exactly twice those of the off-diagonal elements, and there is no difference between the real and imaginary components. The autocorrelation of any component is peaked at zero frequency separation, while the cross correlation of a real component with an imaginary component is zero at zero frequency separation. The negative sign in Eqs. 2 and 4 is a matter of convention for the phase evolution of $Z(f)$ [36]. Because microwave network analyzers utilize a convention in which the phase decreases with increasing frequency, corresponding to the clockwise motion of $Z(f)$ in the $R+iX$ plane seen in Fig. 1, we include the negative sign.
@@
The second statistical characterization we consider, motivated by the path traced by $Z(f)$ in the complex plane as frequency is varied, is how often an element of the impedance matrix comes within a small distance $\epsilon$ of a target value $Z^{\text{t}}$. If the cavity losses are large, such that the $Q$-widths of modes are larger than the frequency spacing between modes, the two point frequency impedance correlation functions decay in a frequency interval $\gamma$. Thus one can estimate that typically over a frequency range of $\gamma$, the impedance value has traced a path of length $R^{\text{RMS}}$. We now imagine that the path in the $R+iX$ plane, rather than being a one-dimensional line, is broadened to a width $\epsilon$. After $N$ frequency intervals of length $\gamma$, the broadened path will cover an area $N\epsilon R^{\text{RMS}}$. That path is confined approximately to be within a circle of radius $R^{\text{RMS}}$. Equating the area covered by the path to the area $\pi(R^{\text{RMS}})^{2}$ within a circle of radius $R^{\text{RMS}}$, we determine the expected number $N$ of frequency intervals of length $\gamma$, needed for the impedance value to come within $\epsilon$ of a target value $Z^{\text{t}}$ within the circle, which is $N=\frac{\pi R^{\text{RMS}}}{\epsilon}$. This gives a simple estimate for the typical frequency interval needed for the impedance to come within $\epsilon$ of a target value as $\Lambda=\frac{\pi\gamma R^{\text{RMS}}}{\epsilon}$.
@@
We conclude that the frequency interval $\Lambda$ scales proportionally with the mean mode spacing $\Delta$ and inversely with the tolerance $\epsilon$. The dependence on loss is more complicated, however, as there is a hidden and non-trivial $\alpha$ dependence in the distribution function $\mathcal{P}(Z)$. Our theory results apply in the over moded limit, where the spacing between modes is much less than the typical $Q$-width of the mode resonances. This corresponds to the RCM loss parameter $\alpha\gg 1$. In this case multiple modes contribute to the response at any frequency, the details of the distribution of mode spacings become unimportant, and the values of the impedance matrix elements become Gaussian random variables. We can approximate the probability ratio in this limit ($\alpha\gg 1$) to take the form
[…]
depending on if $Z$ is a diagonal or off-diagonal element of the RCM normalized impedance matrix.
@@ 3 Experimental Setup and Data Analysis
In this section we introduce the experimental platforms and methods used to verify the theoretical predictions in Sec. 2. Our empirical data comes from three kinds of microwave systems which we characterize by the number of dimensions of wave propagation $\mathcal{D}$. Schematics of the systems are provided in Fig. 3: (b) a tetrahedral graph ($\mathcal{D}=1$) [55, 56, 57, 58], (c) a chaotic quarter bowtie billiard ($\mathcal{D}=2$) [59, 60, 61, 62, 63, 63, 64, 65], and (d) a three dimensional cavity with various geometrical symmetry breaking elements ($\mathcal{D}=3$) [66, 67, 68]. The microwave scattering within these cavities is considered complex when the systems are excited with waves whose wavelengths are small compared to the system size. The interference effects for waves following different ray trajectories are then extremely sensitive to details of the system configuration. We use embedded tunable perturbers to electronically change the system configuration in situ in between measurements. The tunable perturbers are illustrated in green in Fig. 3. In a graph, we use voltage controlled mechanical phase shifters [54] which act as variable-length cables, allowing us to change the interference conditions at the nodes. In the billiard and three dimensional cavity we use globally-biased, voltage controlled varactor-loaded metasurfaces [69, 70] which change the amplitude and phase of reflected waves. In the three dimensional cavity we also have a mechanical mode stirrer attached to the ceiling which can be rotated, substantially altering the wave propagation paths. In reverberant scattering environments, most waves will interact with these tunable scattering elements multiple times, allowing for strong control over the wave scattering properties of each system.
@@
where $Z_{0}$ is a diagonal matrix of the characteristic impedances of the scattering channels ($50~\Omega$, in our case), and $M$ is the number of scattering channels. We then calculate the normalized universal fluctuating impedance $Z$ via the formula [25, 20, 21]
@@
where $\langle Z^{\text{raw}}\rangle$ is a matrix whose elements are the ensemble averages of the elements of $Z^{\text{raw}}$. An estimate for the cavity loss $\alpha$ of the system can then be obtained by simultaneously matching the $2M^{2}$ PDFs of real and imaginary components of the RCM normalized impedance matrix to those predicted by Random Matrix Theory [71, 72]
@@
Experimentally, we made ensembles with different values of $\alpha$ by measuring in different frequency bands, changing the overall size of the systems, and distributing attenuators on the bonds of a graph [58, 54, 32]. More details about the experimental systems used in this paper are discussed in Refs. [39, 37, 38], and an example of determining an experimental ensemble $\alpha$ value through the RCM normalization process is provided in Sec. 6.
@@
Because all diagonal elements $Z_{11},Z_{22},\dots,Z_{MM}$ are statistically identical, only one of these quantities needs to be considered, and the same is true for the off-diagonal elements. This means $M=2$ channels is sufficient to show all the unique statistics in this paper. Note that this is only because we are using the RCM normalized impedance, and is not necessarily true for a raw impedance matrix.

### v2

@@ 1 Introduction
Random Matrix Theory (RMT) based models have emerged as especially successful at predicting certain universal properties of wave chaotic systems [Wigner1947, Lyon1975, Couchman1992, Beenakker1997, Guhr1998, Guhr1989, Alhassid2000, Mendez2003, Fyodorov2004, Kuhl2005, Weidenmuller2009, Mitchell2010, Langley2010, Lawniczak2011, Maksimov2011, Kuhl2016, Li2017, Grabsch2018, Hougne2020]. One such model, the Random Coupling Model (RCM), was developed to incorporate nonuniversal physical features that practical systems invariably have, such as coupling details of ports or antennas [Zheng20051Port, Zheng20052Port] and short orbits (also known as direct processes) [Hart2009, Yeh2010, Lee2013], with an RMT description of the wave propagation within the enclosure. The RCM models the system including its ports in terms of an impedance or admittance matrix that linearly relates the voltages and currents appearing at the system’s ports. From these matrices, combined with knowledge of the external transmission systems feeding the ports, it is possible to evaluate the system scattering matrix, which is often of interest.
The RCM has been applied to understanding the statistical electromagnetic properties of complex enclosed systems for decades [Hemmady2005, GilGil2016, Ma2023b], and has also been empirically verified to describe acoustic scattering [Auregan16]. Predictions of the RCM have been compared to experimental scattering and impedance matrices [Hemmady2005, Hemmady2005b, Yeh2010, Yeh2010b, Lawniczak2010, Lawniczak2019], admittance matrices [Hemmady2006c], conductance fluctuations [Hemmady2006], fading statistics [Yeh2012], complex time delay statistics [Lei2021Stats, Shaibe2025_CTD], and scattering singularity statistics [Shaibe2025_TOP].
One of the applications of the RCM is identifying the underlying universal fluctuations of impedance from experimental ensemble data, which is done by normalizing the measured cavity impedance matrix by the radiation impedance specific to the ports or antennas used in the measurement. The radiation impedance of an antenna characterizes how well it couples to electromagnetic fields in an environment where enclosing boundaries have been removed. This provides specific baseline characterization of the coupling at a port using that antenna. One way to measure the radiation impedance is to put the antenna in an anechoic chamber, equivalent to letting the antenna radiate out to infinity with no reflections, and measuring the resulting scattering parameters. The radiation impedances of the ports are then used to normalize the actual impedance matrix, yielding a dimensionless impedance matrix $Z$ that is independent of many details of the system, allowing for comparison with universal RMT predictions. In this paper, we examine the RCM normalized impedance matrix obtained from experimental data, and identify characteristics of the two point frequency correlation functions of the real (resistance $R$) and imaginary (reactance $X$) components of the impedance matrix elements.
@@
We then apply the knowledge of the correlation functions to address the question: If a port of the system presents a particular value of the complex impedance $Z^{\text{t}}$ at a frequency $f$, what is the average frequency interval $\Lambda$ before the same value of impedance $Z^{\text{t}}$, within some specified tolerance, is seen again? An example is shown in Fig. 1 where we show the path in the complex plane traced by an off-diagonal element of the RCM normalized impedance matrix, $Z_{12}$, over a short frequency range. The arbitrarily chosen target impedance value is $Z_{12}^{\text{t}}=0.082+0.114i$ marked by the black cross, and the small gray circle around it represents a tolerance threshold of $\epsilon=0.015$. The color bar represents the frequency difference $\delta=f_{a}-f_{b}$ (where $f_{b}$ is the first frequency at which $Z_{12}$ is within the gray circle) normalized by a characteristic scale corresponding to the typical $Q$-widths of the modes $\gamma=2\alpha\Delta=0.66~\text{MHz}$. Here $\alpha$ is the dimensionless RCM loss parameter which represents the degree of overlap of the modes ($Q$-width normalized to the mean mode spacing), while $\Delta$ is the mean frequency spacing of the modes.
@@
Empirically we find that the expected frequency interval $\Lambda$ between repetitions of a target impedance depends on the mean mode spacing $\Delta$ of the system, the degree of absorption $\alpha$ of the system, the value of the desired target impedance $Z^{\text{t}}$, and the tolerance $\epsilon$ on the target value. We claim the scaling of $\Lambda$ to be significant for several reasons. $\Lambda$ is a new statistical observable for chaotic scattering that plays a similar role to level-spacing statistics for eigenfrequencies, encoding how frequency correlations determine the recurrence of observable quantities. $\Lambda$ is directly analogous to the Poincaré recurrence time in a non-autonomous, 2 degree of freedom, continuous time dynamical chaotic system, and also parallels the mean first-passage time in stochastic processes. Our scaling law provides a predictive tool for how often a desired impedance condition occurs naturally in frequency for a given cavity or resonator , which has practical relevance for system engineering. If, for example, one wants to use impedance as a sensor, such as of gas concentration or cavity defect formation, by comparing a measured impedance value with a target $Z^{\text{t}}$, then $\Lambda$ would be an important quantity to consider. A large recurrence interval would mean the impedance spectrum is smooth with only a few and distant features, making impedance not useful for such applications due to a lack of sensitivity to perturbations. But with too small of a $\Lambda$, one may encounter issues because the impedance value naturally repeats with even minor perturbations. This is also the first step towards a more difficult question, which would be looking for repetitions not of individual impedance elements, but of special conditions on the entire $Z$ matrix, such as impedance matrix Exceptional Points. Exceptional Point sensors have been proposed which take advantage of the sublinear splitting of the eigenvalues upon perturbation, making them the subject of considerable interest for a variety of sensing applications [Wiersig2014, Zhicheng2019, Chen2025]. Impedance matrix Exceptional Points have previously been shown to be identical to scattering matrix Exceptional Points [Erb2025], so the impedance matrix may serve as a useful platform to study these and other wave scattering singularities.
@@
In this manuscript, we first present the theoretical predictions for the correlation functions and repetition interval $\Lambda$ in Sec. 2. The experimental systems used are introduced in Sec. 3, alongside an explanation of how we apply the Random Coupling Model to the raw data and the statistical methods used in the analysis. The results are shown and interpreted in Sec. 4, and we conclude our work and discuss future research in Sec. 5. In Table 1 of Appendix A we list the relevant parameters that are utilized in this work.
@@ 2 Theory
General forms of two point frequency correlation functions for the fluctuations of the impedance and scattering matrices were first proposed by Fyodorov, Savin, and Sommers [Fyodorov2005, Savin2006]. Various scattering matrix correlation functions have been studied since, often as a way to probe what is known as the elastic enhancement factor [Hemmady2006c, Dietz2009, Dietz2010, Lawniczak2010, Lawniczak2011, Lawniczak2012, Yeh2013, Lawniczak2015, Sokolov2015, Ericson2016, Ramos2016, Bialous2020, Bereczuk2021, Davy2021, Bialous2023, Lawniczak2023]. In this paper, we consider the impedance two point autocorrelation functions, which have not been experimentally investigated before, to our knowledge.
Under the limiting assumptions of ideal coupling (which we achieve through the RCM normalization process), reciprocity (Dyson symmetry class $\beta=1$), and large loss ($\alpha\gg 1$, meaning the modes are strongly overlapping), we arrive at the reduced expressions for the impedance two point frequency correlation functions for the elements of the RCM normalized impedance matrix with condensed notation:
@@
In these equations, $\delta=|f_{a}-f_{b}|$ is the difference in frequency between the two points, $\gamma=2\alpha\Delta$ is the characteristic frequency interval specific to the system, and $R^{\text{RMS}}=\sqrt{\langle(R_{pp^{\prime}})^{2}\rangle}$ is defined as the standard deviation of the fluctuation of the real component of an off-diagonal element of the RCM normalized impedance matrix. Note that in the case of large $\alpha$, $R^{\text{RMS}}$ is known to depend only on $\alpha$ by the formula $R^{\text{RMS}}=(\pi\alpha)^{-1/2}$ [Hemmady2005, Zheng20051Port].
@@
Figure 2: Two point frequency correlation function of the real part of a diagonal impedance element $\langle\tilde{R}(f_{a})\tilde{R}(f_{b})\rangle_{C}$ normalized by the zero separation correlation $\langle(\tilde{R}_{pp})^{2}\rangle$ as a function of $\delta/\gamma=\frac{|f_{a}-f_{b}|}{2\alpha\Delta}$. Comparing prediction of Eq. 1 illustrated by the dashed black curve, and expression derived by Fyodorov, Savin, and Sommers in Ref. [Fyodorov2005], represented by the colored curves corresponding to different choices of loss parameter $\alpha$. The green curve which has $\alpha=10$ is indistinguishable from the dashed black curve.
@@
The correlation functions for the diagonal elements come out to be exactly twice those of the off-diagonal elements, and there is no difference between the real and imaginary components. The autocorrelation of any component is peaked at zero frequency separation, while the cross correlation of a real component with an imaginary component is zero at zero frequency separation. The negative sign in Eqs. 2 and 4 is a matter of convention for the phase evolution of $Z(f)$ [Lei2021Stats]. Because microwave network analyzers utilize a convention in which the phase decreases with increasing frequency, corresponding to the clockwise motion of $Z(f)$ in the $R+iX$ plane seen in Fig. 1, we include the negative sign.
In Fig. 2 we compare Eq. 1 to the correlation function of the real part of the diagonal impedance using Eqs. [10a-b] of Ref. [Fyodorov2005] with $\beta=1$ (reciprocity), all $\kappa=1$, (ideal coupling), and three different $\alpha$ values. We see that for large $\alpha$, our equation and the expression derived by Fyodorov, Savin, and Sommers for the two point frequency correlation function of the real part of the diagonal impedance elements are equivalent. Notice that the correlation function does not change dramatically with $\alpha$, so while our expressions are not perfect they can function as fair approximations even at medium to low loss and weakly overlapping modes. This is true not only in the case of $\langle\tilde{R}_{pp}(f_{a})\tilde{R}_{pp}(f_{b})\rangle_{C}$, but also for the correlation function of the imaginary part of the diagonal impedance $\langle X_{pp}(f_{a})X_{pp}(f_{b})\rangle_{C}$. When it comes to the off-diagonal impedance, the correlation functions are less sensitive to $\alpha$, so Eqs. 3-4 should be even more successful approximations than Eqs. 1-2 at low loss.
The absence of substantial change in the two point frequency correlation functions with $\alpha$ tells us something profound and unexpected. At small $\alpha$, the modes have a narrow $Q$-width relative to the mode spacing. The correlation function over a small frequency band should then look like the correlation function of an isolated mode uninfluenced by any neighbors, which would be a Lorentzian with a scale of the mode $Q$-width $\gamma$. In Fig. 2, we see that the correlation width is nearly the same for overlapping as well as for non-overlapping modes.
@@
The second statistical characterization we consider, motivated by the path traced by $Z(f)$ in the complex plane as frequency is varied, is how often an element of the impedance matrix comes within a small distance $\epsilon$ of a target value $Z^{\text{t}}$. If the cavity losses are large, such that the $Q$-widths of modes $\gamma$ are larger than the frequency spacing between modes $\Delta$, the two point frequency impedance correlation functions decay in a frequency interval $\gamma$. Thus one can estimate that over a frequency range of $\gamma$, the impedance value has traced a path of length $R^{\text{RMS}}$. We now imagine that the path in the $R+iX$ plane, rather than being a one-dimensional line, is broadened to a width $\epsilon$. After $N$ frequency intervals of length $\gamma$, the broadened path will cover an area $N\epsilon R^{\text{RMS}}$. That path is confined approximately to be within a circle of radius $R^{\text{RMS}}$. Equating the area covered by the path to the area $\pi(R^{\text{RMS}})^{2}$ within a circle of radius $R^{\text{RMS}}$, we determine the expected number $N$ of frequency intervals of length $\gamma$, needed for the impedance value to come within $\epsilon$ of a target value $Z^{\text{t}}$ within the circle, which is $N=\frac{\pi R^{\text{RMS}}}{\epsilon}$. This gives a simple estimate for the typical frequency interval needed for the impedance to come within $\epsilon$ of a target value as $\Lambda=\frac{\pi\gamma R^{\text{RMS}}}{\epsilon}$.
@@
We conclude that the frequency interval $\Lambda$ scales proportionally with the mean mode spacing $\Delta$ and inversely with the tolerance $\epsilon$. The dependence on loss is more complicated, however, as there is a hidden and non-trivial $\alpha$ dependence in the distribution function $\mathcal{P}(Z)$. Our theory results apply in the over moded limit, where the spacing between modes $\Delta$ is much less than the typical $Q$-width of the modes $\gamma$. This corresponds to the RCM loss parameter $\alpha\gg 1$. In this case multiple modes contribute to the response at any frequency, the details of the distribution of mode spacings become unimportant, and the values of the impedance matrix elements become Gaussian random variables. We can approximate the probability ratio in this limit ($\alpha\gg 1$) to take the form
$\frac{\mathcal{P}(Z^{\text{peak}})}{\mathcal{P}(Z^{\text{t}})}\approx\mathcal{P}(Z^{\text{peak}})\begin{cases}\exp\left[\frac{(X^{\text{t}})^{2}+(R^{\text{t}}-1)^{2}}{(R^{\text{RMS}})^{2}}\right],&\text{Diagonal}\\ \,\exp\left[\frac{(X^{\text{t}})^{2}+(R^{\text{t}})^{2}}{(R^{\text{RMS}})^{2}}\right],&\text{Off-Diagonal}\end{cases}$ (7)
depending on whether it is a diagonal or off-diagonal element of the impedance matrix that is being considered.
@@ 3 Experimental Setup and Data Analysis
In this section we introduce the experimental platforms and methods used to verify the theoretical predictions in Sec. 2. Our empirical data comes from three kinds of microwave systems which we characterize by the number of dimensions of wave propagation $\mathcal{D}$. Schematics of the systems are provided in Fig. 3: (b) a tetrahedral graph ($\mathcal{D}=1$) [Kottos1997, Kottos2000, Hul2004, Hul2005], (c) a chaotic quarter bowtie billiard ($\mathcal{D}=2$) [Doron1990, Sridhar1991, Graf1992, Alt1998, Stockmann1999, Stockmann1999, Kumar2013, Dietz2015], and (d) a three dimensional cavity with various geometrical symmetry breaking elements ($\mathcal{D}=3$) [Alt1997, Lawniczak2009, Selemani2015]. The microwave scattering within these cavities is considered complex when the systems are excited with waves whose wavelengths are small compared to the system size. The interference effects for waves following different ray trajectories are then extremely sensitive to details of the system configuration. We use embedded tunable perturbers to electronically change the system configuration in situ in between measurements. The tunable perturbers are illustrated in green in Fig. 3. In a graph, we use voltage controlled mechanical phase shifters [Lawniczak2023] which act as variable-length cables, allowing us to change the interference conditions at the nodes. In the billiard and three dimensional cavity we use globally-biased, voltage controlled varactor-loaded metasurfaces [Sleasman2023, Erb2024] which change the amplitude and phase of reflected waves. In the three dimensional cavity we also have a mechanical mode stirrer attached to the ceiling which can be rotated, substantially altering the wave propagation paths. In reverberant scattering environments, most waves will interact with these tunable scattering elements multiple times, allowing for strong control over the wave scattering properties of each system.
@@
where $Z_{0}$ is a diagonal matrix of the characteristic impedances of the scattering channels ($50~\Omega$, in our case), and $M$ is the number of scattering channels. We then calculate the universal fluctuating impedance $Z$ with system-specific features normalized via the formula [Hemmady2005, Zheng20051Port, Zheng20052Port]
@@
where $\langle Z^{\text{raw}}\rangle$ is a matrix whose elements are the ensemble averages of the elements of $Z^{\text{raw}}$. In an ensemble with a high degree of statistical independence between the realizations, we can expect $\langle Z^{\text{raw}}\rangle$ to approach the radiation impedance in the limit of the number of realizations going to infinity, though for practical purposes a few hundred realizations is almost always sufficient. The advantage of using the ensemble average over the radiation impedance in the normalization process is that $\langle Z^{\text{raw}}\rangle$ also contains information about the short orbits of the cavity, which are another source of non-universal effects [Yeh2010, Yeh2010b]. An estimate for the cavity loss $\alpha$ of the system can then be obtained by simultaneously matching the $2M^{2}$ PDFs of real and imaginary components of the RCM normalized impedance matrix to those predicted by Random Matrix Theory [Gradoni2014, Fu2017]
@@
Experimentally, we made ensembles with different values of $\alpha$ by measuring in different frequency bands, changing the overall size of the systems, and distributing attenuators on the bonds of a graph [Hul2005, Lawniczak2023, Lawniczak2019]. More details about the experimental systems used in this paper are discussed in Refs. [Erb2025, Shaibe2025_CTD, Shaibe2025_TOP], and an example of determining an experimental ensemble $\alpha$ value through the RCM normalization process is provided in Appendix B.
@@
Because all diagonal elements $Z_{11},Z_{22},\dots,Z_{MM}$ are statistically identical, only one of these quantities needs to be considered, and the same is true for the off-diagonal elements. This means $M=2$ channels is sufficient to show all the unique statistics in this paper. Note that this is only because we are using the RCM normalized impedance, and is not necessarily true for a raw impedance matrix which would have system-specific features that could be different at each port.